OpenStudy (anonymous):

The circumference of a peg is 3 in. A hole in a piece of wood has a 1 inch diameter. Will the peg go into the hole? Explain, please.

OpenStudy (anonymous):

TT1^2\[\ge\]3 9.869[\ge\]3

OpenStudy (amistre64):

3 = 2pi r 3/2pi = r ; 2r = diam 3/pi = 0.954... which is smaller than 1 inch so the peg is too big